OpenStudy (anonymous):

Simplify: 5(2m + 6n - 3n)

OpenStudy (anonymous):

2methods are there 1) 5(2m+6n-3n) = 5(2m+3n) = 5(2m)+5(3n) == 10m + 15n 2) 5(2m+6n-3n) = 5(2m) + 5(6n) - 5(3n) = 10m+30n-15n=10m+15n answers are same i think 1sst one is shorter just remember to simplify like terms

OpenStudy (anonymous):

10m+30n-15n = 10m+15n